OpenStudy (anonymous):

What is the exact value of the square root of 50 over 3 the square root of 15.? the square root of 30 over 9. 25 the square root of 30 all over 45. 3 the square root of 30 all over 5. 5the square root of 30.

OpenStudy (anonymous):

Do you mean \(\huge \frac{\sqrt{50}}{3\sqrt{15}}\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, well there are some nice square root rules that let us rewrite a problem like this. First rule: \(\Large \sqrt{a}*\sqrt{b} = \sqrt{a*b}\) Second rule: \(\huge \frac{\sqrt{a}}{\sqrt{b}}=\sqrt{\frac{a}{b}}\)

OpenStudy (anonymous):

alright so now what

OpenStudy (anonymous):

We'll use the first rule to split the top up. We want to break it into two things, and we want one of those things to be a perfect square.

OpenStudy (anonymous):

so how do you do that